### v2

# An Attention-Assisted AI Model for Real-Time Underwater Sound Speed Estimation Leveraging Remote Sensing Sea Surface Temperature Data

Abstract
The estimation of underwater sound velocity distribution serves as a critical basis for facilitating effective underwater communication and precise positioning, given that variations in sound velocity influence the path of signal transmission. Conventional techniques for the direct measurement of sound velocity, as well as methods that involve the inversion of sound velocity utilizing acoustic field data, necessitate on–site data collection. This requirement not only places high demands on device deployment, but also presents challenges in achieving real-time estimation of sound velocity distribution. In order to construct a real-time sound velocity field and eliminate the need for underwater onsite data measurement operations, we propose a self-attention embedded multimodal data fusion convolutional neural network (SA-MDF-CNN) for real-time underwater sound speed profile (SSP) estimation. The proposed model seeks to elucidate the inherent relationship between remote sensing sea surface temperature (SST) data, the primary component characteristics of historical SSPs, and their spatial coordinates. This is achieved by employing CNNs and attention mechanisms to extract local and global correlations from the input data, respectively. The ultimate objective is to facilitate a rapid and precise estimation of sound velocity distribution within a specified task area. The comparative analysis demonstrates that the proposed approach achieves superior performance in terms of both accuracy and stability, exhibiting reduced error rates and enhanced resistance to disturbances when benchmarked against existing advanced techniques.
Index Terms:
self-attention, multimodal data fusion, convolutional neural network (CNN), real-time underwater sound speed profile (SSP) estimation, remote sensing sea surface temperature (SST) data.
I Introduction
Underwater sound velocity construction plays an indispensable role in many applications such as underwater positioning, navigation, timing, communication (PNTC) and target recognition due to its decisive factor in the signal propagation mode [1, 2, 3]. Acoustic propagation speed in marine settings exhibits spatial variability due to the combined effects of temperature, salinity, and pressure [4, 5, 6, 7]. Generally, the sound velocity exhibits a stratified variation with depth, so that sound speed profiles (SSPs) are commonly employed to describe the distribution of sound velocity.
The methodology for acquiring SSPs has been a popular research topic for a long period. In traditional ways, SSPs can be directly measured using various instruments, including the conductivity–temperature–depth profiler (CTD), the expendable CTD (XCTD), and the sound velocity profiler (SVP). Alternatively, SSPs can be derived by inversion based on acoustic field data, which encompasses parameters such as signal transmission time or received signal strength [8, 9, 10, 11, 12]. In general, instrument-based measurements can provide precise SSPs with high–depth resolution, but the measurement process is notably time-intensive. For instance, when the instrument is deployed at a regular speed of 50 meters per minute, it takes at least 80 minutes to measure the sound velocity value within a depth range of 2000 meters to complete the deployment and retrieval of the equipment [11]. As a result, it is difficult to provide real-time sound velocity distributions for underwater communication and positioning through instrument-based measurement. Considering that the distribution of sound velocity can affect the distribution of sound field, researchers have proposed a series of methods to invert the sound velocity distribution by using on-site measured sound field data such as signal propagation time in order to accelerate the process of obtaining sound velocity. There are three main frameworks for inversion methods, including matching field processing (MFP) [13], compressive sensing (CS) [14, 15], and deep learning (DL) [4, 16]. The proposed inversion technique enables fast acquisition of sound speed profiles, offering substantial temporal advantages over conventional conductivity-temperature-depth (CTD) sensors and sound velocity profilers (SVPs) in marine data collection. However, because of the interference of noise in sound field measurement, there is a loss in the accuracy of these sound velocity inversion methods. More importantly, it still takes a certain amount of time to measure the sound field data, and new requirements have been put forward for the deployment of expensive sound field measurement equipment. The implementation of SSP inversion techniques is inherently constrained by their dependence on active sonar measurements, thereby limiting their applicability in regions beyond the operational range of underwater acoustic monitoring networks.
Over the past few years, the fast development of underwater automated sensing equipment has provided increasingly rich vertical observation reference data for exploring the distribution of underwater sound velocity. To achieve fast estimation of SSPs and eliminate the need for underwater on-site operations, various methods have emerged that use empirical sound velocity distribution data or remote sensing data. [17, 18, 19] utilize empirical SSPs data to capture the changes in sound velocity from a time series perspective, enabling real-time estimation and prediction of sound velocity distribution in designated ocean areas. These SSP prediction methods have a single type of input data and low complexity of model construction. However, because of the low time resolution of historical samples, its accuracy performance is not good enough in estimating the sound velocity within a small-scale time range. In fact, the variation of sound velocity distribution in a small-scale spatial area is mainly reflected in the shallow water part, as it is more significantly affected by temperature changes. Marine remote sensing technology provides real-time and reliable high spatiotemporal resolution sea surface temperature (SST) data [20], which can offer initial sea surface condition constraints for estimating underwater sound velocity distribution. [21] proposed a self-organizing map (SOM) neural network model that combines empirical sound velocity data and remote sensing SST data, but the model mainly focuses on the local characteristics of sound velocity distribution, which not only fails to capture the influence of SST data on deep ocean sound velocity, but also fails to capture the spatial correlation of sound velocity distribution. As a result, the accuracy performance has not been significantly improved compared to prediction methods.
In order to fast and accurately estimate the SSP of a given task area without on-site underwater data measurement, we propose an interpretable self-attention-assisted multimodal data fusion convolutional neural network (SA-MDF-CNN) model in this paper to deeply capture the intrinsic relationships among empirical SSP data, remote sensing SST data, and spatial coordinates. The core idea is to capture the local feature correlation of regional sound velocity distribution through CNN, and enhance the learning ability of global feature correlation of multimodal data through attention mechanism. To evaluate the effectiveness of the proposed SA-MDF-CNN model, we conducted experiments using Argo data from the Pacific Ocean and measured data from the South China Sea in 2023. Results show that SA-MDF-CNN can achieve lower root mean square error (RMSE) than other spatial SSP construction methods. The contribution of this paper can be summarized as follows:
  • To achieve real-time estimation of SSP without on-site underwater data measurement, we propose an attention-assisted multimodal data fusion convolutional neural network (SA-MDF-CNN) model. We take into account the mutual influence among empirical sound velocity distribution, SST, and spatial position, and the attention module is designed to enhance the global feature learning ability of the model.
  • To evaluate the accuracy performance of the model, we conducted experiments using Argo data from the Pacific region and further validated it through sea trials in the South China Sea in 2023. The RMSE of both experimental results is superior to other spatial SSP constructing methods.
  • To enhance the interpretability of the model, the weights of the model parameters are visualized. As the number of training iterations increases, the model parameter weights gradually focus on the shallow water part. This is reasonable because through the empirical orthogonal function (EOF) decomposition of historical SSPs, it can be seen that the differences in sound velocity distribution are more prominent in shallow water. Therefore, the weight visualization results indicate that the model can effectively capture the characteristics of sound velocity distribution.
The structure of this article is arranged as follows. Section 2 provides a detailed introduction to relevant works. The overall architecture and functional module description of the SA-MDF-CNN model are provided in Section 3. Section 4 gives the experimental findings, and the final section, Section 5, presents the concluding remarks.
II Related works
The distribution of underwater sound velocity plays an essential role in the energy distribution and propagation trajectory of acoustic signals. Therefore, real-time and accurate estimation of sound velocity distribution has significant value for applications based on communication and positioning technologies such as underwater PNTC systems, and target recognition systems.
The traditional way of obtaining sound velocity distribution is usually through direct measurement using shipborne CTD or SVP equipment, which has the advantage of high accuracy [22, 23]. However, it comes with high economic costs and measurement time expenses. With the development of sensing and network technology, many kinds of underwater environmental observation systems have been established to enhance human understanding of the ocean. In 1979, MUNK and Wunsch first put forward the concept of ocean acoustic tomography that using acoustic field observation information to invert the sound velocity distribution [24, 25]. Since then, there have been three mainstream frameworks for sound velocity inversion, namely MFP [13], CS [14, 15], and DL [4, 16]. The MFP framework mainly consists of four steps. Firstly, EOF decomposition is adopted to extract the principal component features of the regional sound velocity distribution. Then, different feature combinations are generated to form candidate SSPs. Next, the sound field distribution is simulated based on ray theory. Finally, the simulated sound field is matched with the measured sound field to determine the optimal estimation of SSP. To accelerate the search speed of the optimal candidate SSP, Tolstoy introduced simulated annealing algorithm in the MFP, which improved the algorithm execution efficiency but resulted in suboptimal solutions [13]. Afterwards, other heuristic algorithms were combined with MFP frameworks for SSP inversion, but they also obtained suboptimal solutions [26, 27]. To further improve the efficiency of inversion algorithm execution, Bianco [15] and Choo [14] proposed SSP inversion framework based on CS, respectively. In the CS framework, the mapping relationship from the sound field distribution to the sound velocity distribution is established through a matrix, which eliminates the search process for matching terms. However, the matrix relationship introduces linear approximation, thus sacrificing inversion accuracy.
In recent years, deep learning theory has developed rapidly and achieved significant results in constructing unknown nonlinear relationships between different data [28, 29]. In addition, long-term underwater environmental observations have accumulated a large amount of data for marine hydrological research, laying a data foundation for the application of deep learning underwater. To address the drawbacks of computational efficiency and accuracy loss in MFP and CS frameworks, we proposed a auto-encoding feature mapping neural network model for SSP inversion in our early work [4], in which the auto-encoder was created to extract deep robust features so as to reduce the impact of acoustic field measurement errors on the accuracy of sound velocity inversion. Considering the insufficient accumulation of historical sound velocity data in some ocean areas, the DL model is prone to over-fitting and reduces accuracy performance. Therefore, we further propose a few-shot meta learning framework to accelerate the convergence speed of the model [16]. However, the above-mentioned methods based on MFP, CS, and DL all rely on real-time measured acoustic field data, which puts high demands on the deployment of underwater observation equipment. Therefore, these methods not only face high equipment economic costs, but also have limited application scope for areas without sonar measurement systems.
Figure 1: SSP Estimation Structure based on SA-MDF-CNN.
Nowadays, how to eliminate the need for underwater on-site data measurement has become a research hotspot in the field of SSP inversion. Based on this requirement, Liu et al. and Lu et al. both proposed a long-short term memory (LSTM) neural network model for SSP prediction [17, 18], which only requires empirical SSP data. However, due to the low temporal resolution of prior data in most marine areas, this prediction method can only describe the overall trend of sound velocity changes and is difficult to obtain high-precision SSP prediction results. In fact, the SSP estimation methods, that rely on a single data modality (sound field data or historical SSP data), are susceptible to data quality issues, such as high sound field measurement noise or low sound velocity sampling time resolution. To improve the robustness of the sound velocity estimation model, researchers have proposed some multimodal data fusion methods for estimating sound velocity that combining data from different sensors and sources to obtain more comprehensive features than a single data source [30, 21]. Yu et al. proposed a radial basis function (RBF) neural network for SSP estimation that mainly using historical temperature, salinity profile data and average SSP data. Nevertheless, the model is not sensitive to varieties of sound velocity, and the estimation results often approach the average SSP profile, making it difficult to accurately describe the spatiotemporal distribution changes of sound velocity. Xu et al. proposed a physics-inspired SOM model for SSP estimation, which introduced remote sensing SST data. However, SOM can only capture the influence of sea surface temperature on surface sound velocity distribution, and lacks the ability to capture large-scale feature correlations. Therefore, the accuracy of sound velocity estimation is difficult to meet practical application requirements.
To obtain real-time and accurate estimation of sound velocity distribution without underwater on-site data measurement, we fully consider the historical sound velocity distribution patterns of different spatial coordinates, as well as the impact of real-time SST changes on the dynamic characteristics of sound velocity distribution, and propose an interpretable SA-MDF-CNN model. In this model, the local correlation of features will be captured through CNN and global correlation of features will be captured through attention mechanism.
III SA-MDF-CNN Structure for SSP Estimation
To realize real-time and accurate estimation of sound velocity distribution, we propose an SSP estimation structure based on SA-MDF-CNN, which is shown in figure 1. In this paper, the ocean area of 5.5N-45.5N and 150.5E-170.5E are selected as the research scope. The remote sensing SST data, latitude and longitude coordinates, and the first three order feature vectors of historical SSPs decomposed by EOFs are first fused to construct the fusion data as the SA-MDF-CNN training data. Then, the multimodal fusion data of 8 grids around the task region is combined to construct the target SSP through the proposed SA-MDF-CNN model. In the following part, we will provide detailed introductions to data sources, data fusion structures, and the composition of neural network model, separately.
III-A Data sources
Remote sensing SST data are provided by the National Oceanic and Atmospheric Administration (NOAA) [31] with a spatial grid resolution of 0.25 and a temporal resolution of 1 day. The SSP data are obtained from the Chinese Observation and Research Station fo Global Ocean Argo System (Hangzhou) [32] with a spatial grid resolution of 1 and a temporal resolution of 1 month. Based on the objective analysis method of gradient-dependent correlation scale optimal interpolation, the 3D grid data of the subsurface layer (5-2000 meters) that covering 179.5W to 179.5E and 89.5S to 89.5N was constructed, and the observation profile was vertically interpolated to 57 standard layers with unequal intervals.
III-B Fusion data construction
Figure 2: The data fusion structure for a single coordinate.
Figure 3: SA-MDF-CNN model.
In the selected ocean region, we grid the entire area into N longitude-scales of 1, M latitude-scales of 1, and vertical depth of H layers. For any coordinate in the grid, the grid remote sensing SST data are expressed as $T^{s}_{n,m}$, and the monthly average SSP data of the historical period are chosen for data fusion, with the depth of the $h$th layer expressed as $D_{n,m}^{h},n\in N,m\in M$.
Within the divided $N*M$ grid ocean area, each sub-grid $\psi$ with a size of $3*3$ forms a set of training data, and each set of training data will be further divided into input data $X$ and output label data $Y$, with a sliding step of 1 for each sub grid, as shown in Figure 1. Specifically, for a sub-grid $\psi$, the fusion data formed by 8 surrounding coordinates will be the input data $X$, and the center SSP of each sub-grid $\psi$ will be taken as the output label $Y$. If the coordinate of the grid center is $L_{Y}=L_{n,m}$, then the coordinates of the input data are represented as:
$\mathcal{L_{X}}=[$ $L_{n-1,m-1},L_{n-1,m},L_{n-1,m+1},L_{n,m-1},L_{n,m+1},$ (1)
$L_{n+1,m-1},L_{n+1,m},L_{n+1,m+1}],\quad n\in N,\quad m\in M.$
For a single coordinate, the data fusion structure is shown in figure 2, which includes the real-time SST data, latitude and longitude coordinates, and the first three EOF feature vectors of the historical SSPs. EOF decomposition, also known as principal component analysis, is a statistical method used to analyze structural features in matrix data and extract major data feature quantities. Through EOF decomposition, the change information of the original sound velocity field can be easily condensed into the first few principal components and their corresponding spatial functions, so the main information of the sound velocity field can be expressed only by the first few eigenvectors.
To obtain the EOF feature vectors, corresponding 5-year (2015 to 2019) historical SSPs from the target ocean area are first selected and then linearly interpolated with a step size of 1 m. This standardized interpolation is used to more clearly describe the differences in sound velocity values at different depths. Suppose there are $J$ SSPs, and the $j$th SSP is expressed as a vector $\boldsymbol{S_{j}}=[s_{j,1},s_{j,2},...,s_{j,h}]^{T},h=1,2,...,H,H=1976$, where $s_{j,h}$ means the sound velocity value at the $h$th depth layer of the $j$th SSP. Since the SSP has been linearly interpolated, the depth of the $h$th depth layer is h m. Then, the matrix formed by all empirical SSPs will be:
$\mathcal{S_{H,J}}=\begin{bmatrix}s_{1,1}&s_{2,1}&\cdots&s_{J,1}\\ s_{1,2}&s_{2,2}&\cdots&s_{J,2}\\ \vdots&\vdots&\ddots&\vdots\\ s_{1,h}&s_{2,h}&\cdots&s_{J,h}\end{bmatrix}.$ (2)
Based on equation (2), the average SSP $\boldsymbol{S_{0}}=[s_{0,1},s_{0,2},...,s_{0,h}]^{T}$ can be calculated by averaging each row of equation (2). Next, a residual matrix can be constructed by subtracting $\boldsymbol{S_{0}}$ from each column of equation (2):
[…]
where $\mathcal{C_{H,H}}$ is a matrix with orders of $H\times H$. Performing EOF decomposition on $\mathcal{C_{H,H}}$ yields eigenvectors and eigenvector coefficients:
[…]
For the $j$th SSP in $\mathcal{S_{H,J}}$, it can be recovered by:
$\boldsymbol{S_{j}}=\boldsymbol{S_{0}}+\sum_{k=1}^{K}\alpha_{k}\boldsymbol{e_{k}},$ (6)
[…]
Finally, by fusing data from 8 coordinate points, input training data can be obtained. For location $L_{n-1,m-1}$, the fused data input $\mathcal{F}^{X}_{n-1,m-1}$ can be expressed as:
$\mathcal{F}^{X}_{n-1,m-1}=[\boldsymbol{T^{S}_{n-1,m-1}},\boldsymbol{L_{n-1,m-1}},\mathcal{E_{H,K}}],$ (8)
where $\boldsymbol{T^{S}_{n-1,m-1}}=[T^{s}_{n-1,m-1,1},T^{s}_{n-1,m-1,2},...,T^{s}_{n-1,m-1,H}]^{T}$ is actually a vector composed of H copies of $T^{s}_{n-1,m-1}$ at location $L_{n-1,m-1}$, and $\boldsymbol{L_{n-1,m-1}}$ is similar that $\boldsymbol{L_{n-1,m-1}}=[L_{n-1,m-1,1},L_{n-1,m-1,2},...,L_{n-1,m-1,H}]^{T}$. The label output data is $\mathcal{S}^{Y}_{n,m}=[s^{Y}_{n,m,1},s^{Y}_{n,m,2},...,s^{Y}_{n,m,H}]^{T}$ at location $L_{n,m}$.
III-C SA-MDF-CNN Model
The network structure of SA-MDF-CNN is shown in Figure 3, mainly including attention module and convolutional network module. The first layer is the input layer, the size of which is 1976*6*8, including 6 types of physical quantities (remote sensing SST, latitude, longitude, and the first 3 order eigenvectors) at 8 coordinates with a total depth of 1976 meters. The second layer is a flattening layer, which flattens the three-dimensional fusion data into one-dimensional sequence data as input for the self-attention layer. The third layer comes the self-attention layer, which converts the fusion data into query, key and value vectors through three linear transforms. Then, the dot product of each query vector and each key vector is calculated to obtain the attention score matrix, which is further converted into attention weights through softmax function. Finally, the value vectors are summed using attention weights to obtain the final output features. These layers are designed to enhance the model’s ability of paying attention to the relationship between SSPs in different ocean areas in the input fusion data, so as to better capture the global dependence and key information between fusion data and the sound velocity distribution in the target ocean area.
The fourth and fifth layers are functional layers designed to reshape one-dimensional sequence data into two-dimensional data for subsequent feature extraction through convolution. The convolution layer is designed to further extract the features from the output data of the attention layer, which is better at capturing small-scale relationships of input features. The pooling layer is used to reduce the spatial dimension of the data, thereby reducing the number of parameters and reduce computational complexity, while retaining important feature information. The final two layers are the fully connected layer and the regression layer. The fully connected layer integrates the local features into the global features and outputs them to the regression layer. The RMSE is designed to be the loss function for calculating the gradient of the model parameters and updating weight parameters of the model through back propagation (BP) algorithm:
$Loss=\sqrt{\frac{\sum_{h=1}^{H}(\hat{s}^{Y}_{n,m,h}-s^{Y}_{n,m,h})^{2}}{H}},$ (9)
where $\hat{s}^{Y}_{n,m,h}$ is the estimated sound velocity value at depth $h$ of location $L_{n,m}$, and the estimated SSP can be expressed as $\hat{\mathcal{S}}^{Y}_{n,m}=[\hat{s}^{Y}_{n,m,1},\hat{s}^{Y}_{n,m,2},...,\hat{s}^{Y}_{n,m,H}]^{T}$. The SSP estimation algorithm based on SA-MDF-CNN is provided in Algorithm 1.
Algorithm 1 The SSP estimation algorithm based on SA-MDF-CNN.
1:Historical average SSP data $\mathcal{S}_{\mathcal{L_{X}}}^{mon}$, grid remote sensing SST $\mathcal{T}^{s}_{\mathcal{L_{X}}}$, position $\mathcal{L_{X}}$;
2:Initialize the parameters $\theta$, learning rate $\gamma$, and set the network parameters according to Table I;
3:for $k=1$ to $MaxEpoch$ do
4:     for $t=1$ to $maxBatchSize$ do
5:         Select a fusion data sample
6:    $\mathcal{F}^{X}_{n-1,m-1}=[\boldsymbol{T^{S}_{n-1,m-1}},\boldsymbol{L_{n-1,m-1}},\mathcal{E_{H,K}}]$;
7:         $\mathcal{S}^{Y}_{n,m}=[s^{Y}_{n,m,1},s^{Y}_{n,m,2},...,s^{Y}_{n,m,H}]^{T}$;
8:         $Loss_{t}=\sqrt{\frac{\sum_{h=1}^{H}(\hat{s}^{Y}_{n,m,h}-s^{Y}_{n,m,h})^{2}}{H}}$;
9:         $\hat{\theta}\leftarrow\text{Adam}(\nabla_{\theta}\frac{1}{BatchSize}\sum Loss_{t},\theta,\gamma)$;
10:     end for
11:end for
12:Complete training and save SA-MDF-CNN model;
13:Estimate and output target SSP $\mathcal{S}_{n,m}^{es}$ at intermediate coordinates $L^{test}_{n,m}$.
III-D Multi-head self-attention module
Figure 4: The principle of the multi-head self-attention mechanism.
The self-attention mechanism is a special type of attention mechanism used to process sequential data such as text, images, or time series. It allows the model to process the sequence in a way that considers the relationship between each element in the sequence and all other elements. The self-attention mechanism calculates weights by evaluating the similarities between different elements in a sequence to determine how much an element is related to other elements. Specifically, a kind of similarity score is defined to describe the similarity between each element in the sequence and other elements. Then, based on these similarity scores, the weight of each element can be calculated, and a weighted representation can be further obtained by multiplying the weight with the corresponding element. In order to better capture the associations between different types of SSP, multiple attention heads are used in this paper as shown in figure 4.
The diagram presented in Figure 4 illustrates the intricate workings of the multi-head self-attention mechanism, which is a fundamental building block in the architecture of transformer models. Three pivotal components are involved in the self-attention mechanism: the query $\mathcal{Q}$, key $\mathcal{K}$, and value $\mathcal{V}$ vectors. These vectors are derived from the input data through linear transformations, as depicted in the upper right section of the diagram where the input is projected into the query, key, and value spaces using learnable weight matrices.
The query vector serves to assess the relevance or correlation with the key vectors. This correlation is quantified and utilized to compute a weighted sum, which amalgamates the contributions of the value vectors. The key vectors are instrumental in determining the attention distribution function, denoted as $a_{i}$ in the diagram, which signifies the importance or weight of each element within the input sequence. In contrast, the value vectors encapsulate the specific features or numeric values associated with each element, playing a crucial role in the aggregation of information.
The mechanism operates by first computing the dot product between the query and each key vector. To prevent the dot products from growing too large and causing numerical instability, especially when the dimensionality of the keys $d_{k}$ is high, each key is normalized by dividing it by $\sqrt{d_{k}}$. This normalization step is shown in the diagram as part of the attention calculation process. Following this, the softmax function is applied to the normalized dot products to generate a set of attention weights. These weights are then used to compute a weighted sum of the value vectors, as represented by the equation:
$\text{Attention}(\mathcal{Q},\mathcal{K},\mathcal{V})=\text{softmax}\left(\frac{\mathcal{Q}\mathcal{K}^{\top}}{\sqrt{d_{k}}}\right)\mathcal{V}$ (10)
This operation is performed for each head in the multi-head self-attention mechanism, as shown in the lower section of the diagram. Each head attends to different parts of the input sequence, allowing the model to capture a diverse range of dependencies and nuances within the data. The outputs from all heads are then concatenated and subjected to a linear transformation, resulting in the final output of the multi-head self-attention mechanism.
Given an input vector $\mathbf{x}\in\mathbb{R}^{dim}$, represented as $\mathcal{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{M}]$, where $dim$ denotes the number of feature dimensions and $M$ represents the number of sequence samples. Upon receiving a query vector $\mathbf{q}$, the self-attention function is defined as:
$\mathrm{att}(\mathcal{X},\mathbf{q})=\sum_{i=1}^{M}a_{i}\mathbf{x}_{i}=\sum_{i=1}^{M}\frac{\exp(\mathrm{s}(\mathbf{k}_{i},\mathbf{q}))}{\sum_{j=1}^{M}\exp(\mathrm{s}(\mathbf{k}_{j},\mathbf{q}))}\mathbf{v}_{i},$ (11)
where $a_{i}$ is the $i$th attention weight, computed by the softmax function over the scaled dot product of the key vector $\mathbf{k}_{i}$ and the query vector $\mathbf{q}$. The function $\mathrm{s}(\mathbf{k}_{i},\mathbf{q})$ represents the similarity score between the key vector $\mathbf{k}_{i}$ and the query vector $\mathbf{q}$, which is typically calculated as the dot product $\mathbf{k}_{i}^{\top}\mathbf{q}$ divided by the square root of the key vector’s dimension $\sqrt{d_{k}}$. The value vector $\mathbf{v}_{i}$ corresponds to the $i$th sample in the input sequence, and the output of the self-attention function is a weighted sum of these value vectors, where the weights are determined by the attention mechanism.
In a standard self-attention module, the multi-head attention mechanism employs multiple query vectors $\mathcal{Q}$ to select multiple sets of information from the input data in parallel. Specifically, each head independently computes the attention weights between the query vectors $\mathcal{Q}$, key vectors $\mathcal{K}$, and value vectors $\mathcal{V}$ to extract different subsets of information. The output of the multi-head self-attention module is calculated by concatenating the outputs of all heads and then applying a linear transformation. The mathematical expression is:
$\mathrm{MultiHead}(\mathcal{Q},\mathcal{K},\mathcal{V})$ $=\mathrm{Concat}(\mathrm{head}_{1},\ldots,\mathrm{head}_{\mathrm{h}})\mathcal{W}^{O},$ (12)
where $\mathrm{head_{i}}=\text{Attention}(\mathcal{QW}_{i}^{Q},\mathcal{KW}_{i}^{K},\mathcal{VW}_{i}^{V})$, $\mathcal{W}_{i}^{Q}\in\mathbb{R}^{d_{k}},\mathcal{W}_{i}^{K}\in\mathbb{R}^{d_{k}},\mathcal{W}_{i}^{V}\in\mathbb{R}^{d_{v}}$, and $\mathcal{W}^{o}\in\mathbb{R}^{hd_{v}}$ is a projection of the parameter matrix.
IV Results and discussions
To test the effectiveness of the proposed SA-MDF-CNN model, the monthly average fusion data from July 2015 to 2019 were selected as input training data and the monthly average SSP from 2020 were set to be the output label data. Experimental results were compared with other widely used methods for constructing spatial sound velocity distribution, including convolutional neural network (CNN), spatial interpolation (SITP) and mean values (MEAN), respectively. The model parameter settings are given in Table I.
TABLE I: Parameter settings of SA-MDF-CNN
Parameter Value
GPU RTX 3090
input size [1976,6,8]
self-attention [8,32]
filter size [2,2]
number of Channels 256
number of filters 256
convolution stride size [1,1]
pool size [2,2]
pooling stride 2
full connected output size 1976
minimum batch size 16
maximum epochs 100
learning rate $10^{-3}$
drop factor of learning rate 0.1
drop period of learning rate 20
IV-A Accuracy Performance
Figure 5: Remote sensing SST data.
To test the accuracy performance of proposed model in estimating SSPs, real-time remote sensing SST data from different locations and depths were fused with EOF feature vectors to construct the sound velocity field. The remote sensing SST data are shown in figure 5, and a visual display of the fusion data is given in figure 6.
Figure 6: Fusion data of remote sensing SST data, latitude and longitude coordinates, and EOF feature vectors.
Figure 7: Comparison of real-time estimated SSP results of different algorithms with depth of 1976 m, where (a) is the SA-MDF-CNN algorithm, (b) is the CNN algorithm, (c) is the spatial interpolation algorithm, and (d) is the mean value method.
TABLE II: Comparison of RMSE at the same depth:1976 meters.
Number Location RMSE of different algorithms (m/s)
SA-MDF-CNN CNN SITP Mean method
1 7.5°N 156.5°E 0.1125 0.1317 0.2456 0.2591
2 7.5°N 151.5°E 0.2599 0.3084 0.3182 0.3377
3 6.5°N 157.5°E 0.1293 0.1402 0.1730 0.1812
4 8.5°N 163.5°E 0.1152 0.1219 0.2775 0.2903
5 11.5°N 165.5°E 0.2090 0.2149 0.2709 0.2842
6 24.5°N 162.5°E 0.0787 0.1115 0.1232 0.1281
7 26.5°N 163.5°E 0.1181 0.1378 0.1293 0.1367
8 27.5°N 153.5°E 0.1039 0.1239 0.1532 0.1621
9 7.5°N 160.5°E 0.1945 0.2050 0.4025 0.4229
Average 0.1468 0.1661 0.2326 0.2447
The real-time SSP estimation results of SA-MDF-CNN algorithm with depths ranging from 0 to 1976 meters within $7^{\circ}N$ to $28^{\circ}N$, $150^{\circ}E$ to $165^{\circ}E$ are compared with those of CNN, SITP and mean values method as shown in figure 7. From the intuitive representation of the sound velocity heat-map, it can be seen that the sound velocity distribution estimated by SA-MDF-CNN and CNN has a significant improvement in accuracy compared to SITP and Mean value methods, and the estimated values of the proposed model are more similar to the sound velocity distribution of real samples. A comprehensive evaluation of the accuracy performance of various algorithms across different locations at depths ranging from 0 to 1976 meters is presented in Table II. Intuitive observation shows that the average RMSE results of SA-MDF-CNN, CNN, SITP, and Mean value methods are 0.1468, 0.1661, 0.2326, and 0.2447, respectively. Notably, the RMSE of SA-MDF-CNN for real-time SSP estimation is approximately 12% lower than that of CNN and approximately 30-40% lower than that of traditional SITP and mean value methods. Figure 8 gives a comparison of sound velocity estimation curves at $7.5^{\circ}N,156.5^{\circ}E$. In figure 8 (a), the curve of SA-MDF-CNN is the closest to the real SSP, and in figure 8 (b), the error mainly manifests in the shallow ocean area within 500 meters because the sound velocity in shallow water is more severely and irregularly affected by temperature. However, among all methods the error disturbance of SA-MDF-CNN is the smallest.
Figure 8: A comparison example of real-time estimated SSP results of different algorithms with depths ranging from 0 to 1976 meters, where (a) is the comparison of estimated results and real SSP, and (b) is the comparison of error distribution between the estimated results of different algorithms and the real SSP.
Figure 9: Comparison of estimation results of different algorithms at different shallow ocean depths.
TABLE III: Comparison of RMSE at different depth.
Depth Location RMSE of different algorithms (m/s)
SA-MDF-CNN CNN SITP Mean method
200m 6.5°N 159.5°E 0.6075 0.6911 1.085 1.143
7.5°N 156.5°E 0.2911 0.3468 0.6688 0.7035
7.5°N 161.5°E 0.3859 0.4092 1.019 1.085
9.5°N 164.5°E 0.1383 0.1775 0.3831 0.4018
13.5°N 169.5°E 0.3115 0.3556 0.5565 0.5862
Average 0.3469 0.3960 0.7425 0.7839
300m 11.5°N 165.5°E 0.4957 0.5155 0.6903 0.7240
12.5°N 152.5°E 0.1901 0.3242 0.3392 0.3542
13.5°N 154.5°E 0.2889 0.3028 0.4363 0.4594
14.5°N 158.5°E 0.1360 0.1411 0.3860 0.4088
14.5°N 166.5°E 0.2943 0.3446 0.5869 0.6156
Average 0.2810 0.3256 0.4877 0.5124
500m 7.5°N 151.5°E 0.2293 0.2539 0.4839 0.5101
6.5°N 157.5°E 0.2299 0.2613 0.3371 0.353
8.5°N 164.5°E 0.2943 0.3257 0.6033 0.6368
11.5°N 156.5°E 0.1571 0.1625 0.3331 0.3481
17.5°N 163.5°E 0.1133 0.1841 0.2544 0.2721
Average 0.2048 0.2375 0.4024 0.4240
Figure 10: Comparison of error distributions of real-time estimated SSPs for different algorithms at different shallow sea depths.
To further test the accuracy and performance of the model in estimating the sound velocity distribution in shallow water areas, the real-time SSP estimation results of SA-MDF-CNN algorithm at 200 m, 300 m and 500 m depths are compared with those of CNN, SITP and mean values method as shown in figure 9. As can be seen from the enlarged part of each sub-figure, the estimated SSP curves by SA-MDF-CNN at 200 m, 300 m and 500 m depths are significantly closer to the real SSP prediction results than those of CNN, SITP and mean method at the same ocean area. Specifically, an in-depth investigation into the performance of different algorithms in shallow ocean areas at varying depths and locations was conducted, with the real-time SSP estimation RMSE results compared and displayed in Table LABEL:table3. The largest RMSE is observed at a depth of 200 meters, which can be attributed to the increased interference from real-time sea conditions in the shallow sea environment, such as wind speed, wave height, tsunami, and other factors. Concurrently, the experimental results indicate that the estimation accuracy of the proposed SA-MDF-CNN model at 200 meters depth is enhanced by approximately 13% compared to CNN and by approximately 53% compared to SITP and mean value method. At an ocean depth of 300 meters, the proposed algorithm demonstrates an improvement of about 14% over CNN and about 42% over SITP and mean value method. At 500 meters depth, the proposed model shows an improvement of about 14% compared to CNN and about 49% compared to SITP and mean value method. These experimental results collectively illustrate that the proposed model maintains significant precision advantages even in shallow sea environments of varying depths, thereby underscoring its applicability and effectiveness. The results indicate that SA-MDF-CNN is also applicable even in shallow ocean environment with different depths, and still has smaller RMSE compared with other algorithms, which means that the model still has good feature capture ability in shallow ocean scenes with significant changes in sound velocity. The average absolute error distribution comparison of real-time estimated SSP of different algorithms tested in the set of 200m and 500m depths is given in Figure 10. The error fluctuation of SA-MDF-CNN is the most stable, which indicates that SA-MDF-CNN still has better robust performance under shallow sea conditions with different depths.
Figure 11: Interpretability analysis of the SA-MDF-CNN model, where (a) is the first three order EOF the historical SSP at different positions (b) is the attention weight with Epoch=10, (c) is the attention weight with Epoch=50, and (d) is the attention weight at Epoch=10.
IV-B Interpretability analysis
To investigate why the proposed SA-MDF-CNN model has better performance compared to CNN, we conducted an experiment and visually represented the attention weights of the model as shown in figure 11. Figure 11 (a) illustrates the first three order EOF feature vectors of the historical SSPs at various latitude and longitude coordinates. Figure 11 (b), (c), and (d) correspond to the attention weights when the training epochs are 10, 50, and 100, respectively. Through observing the evolution trend of attention focus with increasing training times, we found that the SA-MDF-CNN model increasingly focuses on the sound velocity distribution in shallow waters within 1000 meters. This trend aligns with the first three orders of EOF feature distribution shown in panel (a), although the EOF itself emphasizes the SSP distribution between 0-500 meters. The attention depth of the proposed model spans 0-1500 meters, which is attributed to the variation of SSP distribution under different latitude and longitude coordinates. Therefore, in order to enhance the model’s ability to focus on the global SSP distribution, we added a multi-head self-attention layer after input data fusion. This mechanism further confirms the effectiveness of the proposed model. Experiments based on the aforementioned public dataset demonstrate that the proposed SA-MDF-CNN model outperforms other algorithms in estimating sound velocity in both shallow and deep waters.
IV-C Efficiency and stability
TABLE IV: Comparison of Neural Networks
Network Number of Parameter Train time
CNN 61.9M 72s
SA-MDF-CNN 149M 40s
Although the accuracy performance in SSP estimation of SA-MDF-CNN is better than traditional methods, the disadvantage is that SA-MDF-CNN increases the number of parameters and computational overhead. Table IV shows the comparison of parameters and running time between the proposed model and CNN. It can be seen that the proposed model in this paper needs more parameters and training time than CNN, which is a necessary overhead for performance improvement.
Figure 12: Comparison of RMSE convergence of different algorithms.
To test the stability of the model, the convergence of SA-MDF-CNN and CNN with different numbers of iteration is given in figure 12. Both SA-MDF-CNN and CNN can converge in less than 200 training iterations. The convergence of SA-MDF-CNN is smoother than that of CNN, which may be due to the multi-head self-attention mechanism focusing more on global features and therefore not causing significant parameter updates due to drastic changes in a small range of the data.
V Ocean experiments
Figure 13: Data sampling locations of ocean experiments.
Figure 14: Comparison of real-time SSP estimation results of different algorithms on ocean experimental data, where (a) shows the SSP curves, and (b) shows the error distributions.
In order to verify the practicality of the SA-MDF-CNN model, a comprehensive deep-ocean experiment was conducted in the South China Sea in April 2023. As illustrated in figure 13 (a), various instruments such as CTD and XCTD were employed at diverse marine locations, yielding a total of 14 distinct samples. The depth of sampling reached 3500 meters, with intervals set at every 100 meters. It is evident that the selected location for this experiment does not conform to the rigid 3x3 grid pattern, thus allowing for a more rigorous evaluation of the proposed model. The fusion data at the blue coordinates in figure 13 (b) were selected as the input of the model, and the real-time SSP under the red pentagram was used as the prediction output of the model.
Figure 15: Comparison of real-time estimation SSP results of different algorithms in shallow sea environment with different depths in ocean experiment, where (a) and (c) are SSP curves of 300 meters depth, (b) and (d) are error distributions of 500 meters depth.
Figure 16: RMSE comparison of real-time estimated SSPs by different algorithms of ocean experiments.
To evaluate the effectiveness of the model in estimating the distribution of sound velocity at full-ocean depth, the estimated SSP of SA-MDF-CNN is compared with other algorithms in figure 14. It is evident that the proposed model exhibits lower RMSE and more stable error disturbance compared to traditional algorithms and CNN. To further test the sound velocity estimation performance in shallow waters with significant changes in sound velocity, estimation of sound velocity at 300, 500 depth meters by SA-MDF-CNN are compared with other methods in figure 15. The results indicate that the proposed algorithm has smaller error disturbances and maintains significant accuracy advantages compared to other algorithms in shallow water environments. Figure 16 gives a more detailed comparison of the RMSE results of different algorithms, clearly showing that the proposed model has significant advantages at different depths.
VI Conclusion
To construct a real-time sound velocity field and eliminate the need for underwater onsite data measurement operations, a SA-MDF-CNN model driven by multimodal data fusion is proposed in this paper, which fuses historical SSP features, real-time remote sensing SST, latitude and longitude coordinates. The fusion data is constructed by gridding, and the real-time SSP is retrieved based on the real-time remote sensing SST. The effectiveness of our proposed model was tested in the case of non-standardized grid through ocean experiments. The experimental results show that our proposed model is not only suitable for real-time SSP estimation under non-strictly standardized grid conditions, but also suitable for shallow sea conditions with different depths.
Acknowledgments
The authors acknowledge the historical SSP data support from the China Argo Real-time Data Center, (https://www.argo.org.cn/, latest access:December 20, 2024). The remote sensing SST data can be acquired at the National Oceanic and Atmospheric Administration (https://www.commerce.gov/, latest access: December 25, 2024).